# Leveraging data-driven weather models for improving numerical weather prediction skill through large-scale spectral nudging

### v1

@@ 6.1 Verification with global power spectra
The power spectra of the GDPS-SN forecast variables remain approximately the same as in the GDPS-CTL (Fig. 7), which essentially demonstrates that the new hybrid system can adequately resolve the full range of scales present in the GDPS-CTL, without significant smoothing. The resulting spectral amplitude ratio of transient-eddy anomalies, $\gamma$ (Eq. 1), also has values very close to one for all prognostic variables, lead times and spatial scales for both GDPS-CTL and GDPS-SN. This provides a solid proof-of-concept for the spectral nudging approach (Fig. 8). The only exception is a 10-percent reduction of $\gamma$ for scales between 2000 and 4000 km at lead times approaching day 10 (wavenumbers 10-20 in Figs. 8a, b). This reduction is explained by GC having some variance deficiency at the corresponding range of scales (2000-2750 km) for longer lead times (beyond day 4). Sensitivity tests with time-varying cut-off length scales that more strictly follow GC’s time-evolving effective resolution, showed that this issue could be avoided, but at the expense of significantly reduced GDPS-SN forecast improvements with respect to the GDPS-CTL. This is caused by a weaker constraint being imposed on GDPS-SN towards GC at synoptic scales by such a configuration.
Comparison of the spectral coherence (Eq. 2) between the forecasts and GDPS analysis shows an improved skill of GDPS-SN over GDPS-CTL (Figs. 8c, d). These improvements are somewhat weaker than those seen with GC at shorter lead times (up to day 5), but become comparable beyond day 5 (Figs 2c, d). The scales relevant for improvements at shorter lead times are smaller than those included in the spectral nudging configuration. Nudging the smaller scales would therefore bring the coherence on par with GC but it would lead to variance deficiency, as discussed earlier.
Figure 7: Same as in Fig. 1, but for the analysis, GDPS-CTL and GDPS-SN.
Figure 8: Same as in Fig. 2, but for GDPS-SN and GDPS-CTL.
An illustrative example of the value added by the new hybrid system over GC is provided in Fig. 9. Large-scale features of transient-eddy anomalies in GDPS-SN tend to be better (but not fully) aligned with GC than with GDPS-CTL, while containing similar level of fine-scale details as in GDPS-CTL. Figure 10 suggests that large-scale components of GDPS-SN forecast fields are preconditioned with the piloting AI model. Therefore, further improvements to its predictions have potential for similarly improving GDPS-SN.
Figure 9: 10-day forecast of 700-hPa transient-eddy specific humidity anomalies, valid 22 February 2022 at 12z for (a) GDPS-CTL, (b) GC and (c) GDPS-SN.
@@ 6.2 Verification against radiosondes
Comparison of forecasts at full resolution against the global radiosonde observations clearly indicates that GDPS-SN forecasts are closer to the observations than GDPS-CTL forecasts for basically all variables, lead times and levels, and over both seasons (Fig. 10). The improvements gradually grow in the first 72 h and are remarkably constant with altitude, which contrasts with the strong vertical variations observed with GC in Section 4. This demonstrates that large-scale spectral nudging, despite being only applied in the mid-level layer for winds and temperature, is sufficient to impart similar improvements at both upper- and lower-level layers that are not constrained by spectral nudging. Compared to boreal winter (Fig. 10a), the improvements from GDPS-SN are reduced by a factor of about 2 during boreal summer (Fig. 10b), which is not surprising given that the benefits from GC over GDPS-CTL at the large scales are smaller by about the same factor for zonal wind and temperature in the mid-level (not shown).
In terms of region-wise performances, the GDPS-SN has smaller RMSE than GDPS-CTL in every subdomain except over Asia in the mid- and upper-levels during boreal summer (Figs. 11c and d), which is relatively consistent with the degradations from GC over the GDPS-CTL at the large scales previously noted for the mid-levels over that part of globe during this season (see Fig. 5d). Unlike GC, the improvements from GDPS-SN over GDPS-CTL are virtually identical when only the large-scale components of the forecast are compared to the radiosonde observations (compare the left and right columns in Fig. 11 and, again, in Fig. 6). This does not necessarily imply that the improvements only originate from the large scales. It rather implies that the relative RMSE reductions with GDPS-SN are of similar level for both the large and small scales. In other words, as the small scales are generally conditioned by the large scales, these results suggest that the fine scales predicted by GDPS-SN benefit from fractional improvements that are on par with its large-scale skill enhancement.
Figure 10: Same as in Fig. 4, except that upward-pointing red triangles indicate a reduction of the RMSE in GDPS-SN with respect to the GDPS-CTL, whereas downward-pointing blue triangles indicate the opposite. Only verification using forecasts at full resolution is shown.
@@
It is also interesting to compare GDPS-SN to GC. For the sake of fair comparison, this will only be done at low resolution. Predictions from GDPS-SN are closer to the radiosonde observations than GC in the upper and mid-level layers, except for dewpoint depression. Conversely, for the low-level layer, GC is closer to the observations, except for geopotential height (Fig. 12). The degradations with respect to dewpoint depression, also observed in the mid-level layer where spectral nudging is active, suggest that further gains could likely be achieved by nudging the humidity field in GEM towards GC. The results also suggest that extending spectral nudging to the boundary layer could be beneficial. This, however, would necessitate an AI model with more vertical levels in the boundary layer and fine-tuned to emulate the GDPS analyses rather than ECMWF-ERA5, for improved consistency with the GEM model’s lower-boundary forcings.
The small but statistically significant GDPS-SN improvements over GC in the mid-level layer for temperature and zonal winds, two variables that are nudged toward GC, that grow with the integration time, are somewhat surprising (Fig. 12). This could be due to a positive feedback mechanism between the large and small scales in GEM. The averaging effect caused by the blending of two forecasts due to a long relaxation time, as previously discussed in Section 5, may also be partly responsible for these improvements. Answering this question is, however, deemed outside the scope of this paper.
Finally, the overall region-wise comparison of GDPS-SN and GC forecasts at low resolution (Fig. 13) shows that the results that were observed over the globe hold also relatively well for each of the subdomains. The fact that spectral nudging is applied only in the mid-level layer targeting only the synoptic scales, it considerably limits the ability of GDPS-SN to fully leverage the strengths of GC for seasonal periods and over regions where mesoscale near-surface conditions play a major role in determining the weather. This may explain the degradations seen over the tropics.
@@ 6.3 Verification against ECMWF analyses
The comparison of GDPS-SN and GDPS-CTL forecasts against ECMWF-HRES analyses leads to very much similar results as the comparison against radiosonde observations, with general improvements for all variables, lead times and pressure levels (not shown). Figure 14 provides an example of improvements in anomaly correlation coefficient (ACC). For 500-hPa geopotential heights, the ACC improvements over the Northern Hemisphere region at forecast day 7 are equivalent to an 18-h increase in predictability during boreal winter and to an 8-h increase in summer. The corresponding ACC improvements over the Southern Hemisphere region are about 2 hours during austral winter and even exceed 24 hours during austral summer. There are also notable improvements in zonal wind ACC in the tropics that are equivalent to 24 hours or more (e.g., see Fig. 14b,e). The strength of GC in improving the ACC around day 5 (and beyond) is thus well leveraged by the hybrid system.
Figure 14: Anomaly correlation coefficient against the ECMWF-HRES analyses with GDPS-CTL (blue) and GDPS-SN (red) for (a,d) northern hemisphere 500-hPa geopotential height, (b,e) tropical 850-hPa zonal wind, and (c,f) southern hemisphere 500-hPa geopotential height. Results are presented for 60 cases of boreal winter 2022 (left column: a–c) and 62 cases of boreal summer 2022 (right column: d–f). No marker at a pressure level implies that the null hypothesis, stating that the statistics of the two samples are the same, cannot be rejected based on the 95th percentile.
@@ 6.4 Verification against surface observations
For verification at the surface level, combined SYNOP, METAR and Surface Weather and marine OBservations (SWOB; available only over Canada and distributed by ECCC) data were used for surface pressure, screen-level temperature, dewpoint depression and anemometer-level wind speed. For precipitation, 24-h accumulation ground observations are subject to collection and quality control by the CAnadian Precipitation Analysis system (CaPA; Lespinas et al. (2015)). Any data pertaining to stations with an altitude difference larger than 100 m with respect to GDPS orography, along with any observed wind speed below 1.5 $\text{m s}^{-1}$, are excluded.
@@
Figure 16: Q-Q plots of day-5 GC (black), GDPS-CTL (blue) and GDPS-SN (red) forecasts against combined SYNOP, SWOB and METAR for: (a, b) anemometer-level wind speeds and (c, d) screen-level temperatures over North America for (a, c) 30 cases of boreal winter valid at (a) 00 and (c) 12 UTC, and (b, d) 31 cases of boreal summer valid 00 UTC. Local standard time over North America is mostly between UTC-5 and UTC-8 h. Diamond symbols denote selected percentiles with their values printed. Shadings denote 95% confidence bands for the inverse cumulative distribution functions, based on the Kolmogorov-Smirnov statistic.
@@
The ability of the GPDS-SN to represent high-impact weather, such as high wind speeds and extreme temperatures, is not significantly different from the GDPS-CTL but, at the same time, it is substantially better than GC, which provides a strong example of the value added by the hybrid system over GC (Fig. 16). For example, the distribution of late-afternoon wind speeds over North America provided by GC shows a considerable shift toward lower values, which increases at high percentiles of the distribution, resulting in a -4 $\text{m s}^{-1}$ bias at the 99.9th percentile (Figs. 16 a, b). GDPS-CTL also has a negative bias but affecting only higher percentiles and, more importantly, GDPS-SN closely follows GDPS-CTL, with little or no sign of deceleration at extreme wind speeds for both seasons. GraphCast also suffers a 2-4 K cold bias at mid-to-high temperature percentiles over North America in summer, which has no impact on the GDPS-SN; its distribution aligns very well with the observations. GraphCast’s poor performance with respect to the extremes can be partly explained by its lack of well-resolved fine scales that are generally critical for representing the extremes. Furthermore, inconsistencies in the surface forcing between ERA5 and GDPS analyses may also be a contributing factor for the observed underestimation of temperature extremes by GC. As the spectral nudging configuration presented in this study only targets synoptic scales and the boundary layer is allowed to evolve freely, GDPS-SN is not adversely affected by GC’s limitations regarding the extremes.
@@
Spectral nudging also significantly improves boreal winter precipitation forecast guidance over North America for a large range of thresholds from 1 to 50 mm day-1, mostly by reducing the false alarm ratio, with no reduction of probability of detection, which results in a reduced moist frequency bias and an improved equitable threat score (Fig. 17). These improvements seem to be mostly generated via an improved prediction of synoptic-scale flow. There is, however, not much of an impact of spectral nudging on summer precipitation over North America. Improving precipitation in summer, when weak synoptic forcings are prevalent, may necessitate the application of spectral nudging in the boundary layer.
@@
Figure 17: Quantitative precipitation forecast verification for day 5 assessed by 24-hr precipitation threshold (accumulations between 108 and 132 h of integration) with the (a) equitable threat score, (b) frequency bias, (c) probability of detection, and (d) false alarm ratio. Results from the 60 GDPS-CTL (blue) and GDPS-SN (red) cases of boreal winter are compared to ground observations over North America used by CaPA. Red and blue line markers denote statistically significant differences in favor of GDPS-SN and GDPS-CTL, respectively. Significance is computed by bootstrapping 3-day data blocks consisting of forecast-observation pairs from all stations of a given region. No marker at a threshold level implies that the null hypothesis, stating that the statistics of the two samples are the same, cannot be rejected based on the 90th percentile.
@@ 6.5 Tropical cyclone evaluation
Spectral nudging improves TC trajectories by reducing the errors resulting from GDPS-CTL’s tendency of predicting storms that move too slowly and veer too much to the right (with respect to its observed track), especially at lead times beyond day 5 (Figs. 18a and b). Conversely, there is little or no significant impact of spectral nudging on the TC intensity, as measured by the minimum central pressure and maximum sustained wind speed (Figs. 18c and d). Both GDPS-CTL and GDPS-SN have a similar week intensity bias, which is a well-known weakness of the GEM model (see McTaggart-Cowan et al. (2024), for more details). Severe TC weak-intensity biases are typical of state-of-the-art deterministic AI-based forecast models (e.g., Bouallègue et al. (2024)) due to the smoothing effect caused by learning to minimize the MSE. However, AI-based models, including GC, are generally capable of predicting the TC trajectories with enhanced accuracy (Lam et al., 2022). Therefore, the improvements in the predicted trajectories by GDPS-SN show that the proposed hybrid system is capable of leveraging the well-documented strength of the AI-based models, without sacrificing TC intensity. Furthermore, these results suggest that the weak-intensity bias for TCs could be reduced in future by focusing on improving the physics-based NWP component of the hybrid system.
Figure 18: Tropical cyclone tracking performance comparison of the GDPS-CTL (blue) and GDPS-SN (red) against IBTrACS data in three northern hemisphere basins (West Atlantic, East and West Pacific) for 62 cases of boreal summer 2022: (a) along-track and (b) cross-track mean position errors, (c) mean central pressure error and (d) mean maximum wind speed error. Negative along-track errors imply cyclones move too slowly and positive cross-track errors imply cyclones veer to much to the right with respect to the observed trajectory.

### v2

@@ 6.1 Verification with global power spectra
The power spectra of the GDPS-SN forecast variables is indistinguishable from GDPS-CTL (not shown), which essentially suggests that the new hybrid system can adequately resolve the full range of scales present in the GDPS-CTL, without significant smoothing. The resulting spectral amplitude ratio of transient-eddy anomalies, $\gamma$ (Eq. 1), also has values very close to one for all prognostic variables, lead times and spatial scales for both GDPS-CTL and GDPS-SN (Fig. 7). This provides a solid proof-of-concept for the spectral nudging approach. The only exception is a 10-percent reduction of $\gamma$ for scales between 2000 and 4000 km at lead times approaching day 10 (wavenumbers 10–20 in Figs. 7a, b). This reduction is explained by GC having some variance deficiency at the corresponding range of scales (2000–2750 km) for longer lead times (beyond day 4). Sensitivity tests with time-varying cut-off length scales that more strictly follow GC’s time-evolving effective resolution showed that this issue could be avoided, but at the expense of significantly reduced GDPS-SN forecast improvements with respect to the GDPS-CTL. This is caused by a weaker constraint being imposed on GDPS-SN towards GC at synoptic scales by such a configuration.
Comparison of the spectral coherence (Eq. 2) between the forecasts and GDPS analysis shows an improved skill of GDPS-SN over GDPS-CTL (Figs. 7c, d). These improvements are somewhat weaker than those seen with GC at shorter lead times (up to day 5), but become comparable beyond day 5 (Figs 2c, d). The scales relevant for improvements at shorter lead times are smaller than those included in the spectral nudging configuration. Nudging the smaller scales would therefore bring the coherence on par with GC, particularly for short lead times, but it would result in variance deficiency, as discussed earlier.
Figure 7: Same as in Fig. 2, but for GDPS-SN and GDPS-CTL. Additionally, for convenience of comparison, GC has been added in Fig. 7c and d.
An illustrative example of the value added by the new hybrid system over GC is provided in Fig. 8. For temperature, which is explicitly nudged in GDPS-SN, the large-scale features of its transient-eddy anomalies tend to very closely resemble GC (Fig. 8a–c). Even for specific humidity, a variable which is not directly constrained by nudging, the large scales are also found to be more aligned with GC than with GDPS-CTL (Fig. 8d–f). This is however not surprising as it reinforces the findings of previous studies on spectral nudging (Husain et al., 2014). Regarding fine scales, Fig. 8 clearly shows that GDPS-SN has similar level of variability as GDPS-CTL, which is not the case with GC as it clearly suffers from severe smoothing.
Figure 8: 10-day forecast of 700-hPa transient-eddy anomalies for (a–c) temperature and (d–f) specific humidity, valid 22 February 2022 at 12z for GDPS-CTL (left), GC (middle) and GDPS-SN (right).
@@ 6.2 Verification against radiosondes
Comparison of forecasts at full resolution against the global radiosonde observations clearly indicates that GDPS-SN forecasts are closer to the observations than GDPS-CTL forecasts for basically all variables, lead times and levels, and over both seasons (Fig. 9). The improvements gradually grow in the first 72 hr and are remarkably constant with altitude, which contrasts with the strong vertical variations observed with GC in Section 4. This demonstrates that large-scale spectral nudging, despite being only applied in the mid-level layer for winds and temperature, is sufficient to impart similar improvements at both upper- and lower-level layers that are not directly constrained by spectral nudging. Compared to boreal winter (Fig. 9a), the improvements from GDPS-SN are reduced by a factor of about 2 during boreal summer (Fig. 9b), which is not surprising given that the large-scale improvements with GC over GDPS-CTL are smaller by about the same factor during boreal summer (not shown).
In terms of region-wise performances, the GDPS-SN has smaller RMSE than GDPS-CTL in every subdomain except over Asia in the mid- and upper-levels during boreal summer (Figs. 10c and d), which is consistent with the degradations from GC over the GDPS-CTL at the large scales in this region (Fig. 5d). Unlike GC, the improvements from GDPS-SN over GDPS-CTL are virtually identical when only the large-scale components of the forecast are compared to the radiosonde observations (compare the left and right columns in Fig. 10 and, again, in Fig. 5). This does not necessarily imply that the improvements only originate from the large scales. It rather implies that the relative RMSE reductions with GDPS-SN are of similar level for both the large and small scales. In other words, as the small scales are generally conditioned by the large scales, these results suggest that the fine scales predicted by GDPS-SN benefit from fractional improvements that are on par with its large-scale skill enhancement.
Figure 9: Same as in Fig. 4, except that upward-pointing red triangles indicate a reduction of the RMSE in GDPS-SN with respect to GDPS-CTL, whereas downward-pointing blue triangles indicate the opposite. Only verification using forecasts at full resolution is shown.
@@
It is also interesting to compare GDPS-SN to GC. For the sake of fair comparison, this is only done at low resolution. Predictions from GDPS-SN are closer to the radiosonde observations than GC in the high- and mid-level layers, except for dewpoint depression. Conversely, for the low-level layer, GC is closer to the observations, except for geopotential height (Fig. 11). The degradations with respect to dewpoint depression, also observed in the mid-level layer where spectral nudging is active, suggest that further gains could likely be achieved by nudging the humidity field in GEM towards GC. Nudging of specific humidity, however, would only be viable once a fine-tuned version of GC that emulates GDPS analyses instead of ERA5 (and ECMWF-HRES) is available. This is expected to bring the mean moisture state of GC inferences closer to GDPS analyses and thereby reduce any error associated with potential adverse response from GEM physics. The results also suggest that extending spectral nudging to the boundary layer could be beneficial. This, however, would necessitate an AI model with more vertical levels in the boundary layer (as mentioned earlier) and fine-tuned to emulate the GDPS analyses rather than ECMWF-ERA5, for improved consistency with the GEM model’s lower-boundary forcings.
The small but statistically significant GDPS-SN improvements over GC in the mid-level layer for temperature and zonal winds, two variables that are nudged toward GC, that grow with the integration time, are somewhat surprising (Fig. 11). The use of a nudging relaxation time $\tau=12$ hr implies that the large-scales in GDPS-SN are not too tightly constrained, which is visible in Fig. 8. This likely allows for some large-scale improvements through upscale propagation of information from an improved representation of the fine scales. However, this choice of $\tau$ also leads to a small averaging effect caused by the blending of two forecasts, as previously discussed in Section 5. To a lesser extent, this may also have contributed to these improvements.
Finally, the overall region-wise comparison of GDPS-SN and GC forecasts at low resolution (Fig. 12) shows that the skill of GDPS-SN is either slightly better than or comparable to GC, with the tropical boundary layer being an exception. Fig. 5 reveals that, compared to GC, GDPS-CTL large scales have substantially reduced skill in the tropics, which is slightly improved with GDPS-SN (see Fig. 10). To leverage the full potential of GC in the tropics it appears to be necessary to extend nudging in the boundary layer, which is not viable at this stage for reasons explained earlier in this section.
@@ 6.3 Verification against ECMWF analyses
The comparison of GDPS-SN and GDPS-CTL forecasts against ECMWF-HRES analyses leads to very similar results as the comparison against radiosonde observations, with general improvements for all variables, lead times and pressure levels (not shown). Figure 13 provides an example of improvements in anomaly correlation coefficient (ACC). For 500-hPa geopotential heights, the ACC improvements over the Northern Hemisphere region at forecast day 7 are equivalent to an 13-hr increase in predictability during boreal winter and to an 8-hr increase in summer. The corresponding ACC improvements over the Southern Hemisphere region are about 3 hours during austral winter and 22 hours during austral summer. There are also notable improvements in zonal wind ACC in the tropics that are equivalent to 30 hours or more (e.g., see Fig. 13b,e). The strength of GC in improving the ACC around day 5 (and beyond) is thus well leveraged by the hybrid system. Note that GC has not been included in this figure as its ACC is strongly influenced by the excessive fine-scale smoothing. To fairly compare GC against the two versions of GDPS, it is necessary to do that at low resolution (as in section 4b). However, based on spectral coherence presented in Fig. 7c–d, low-resolution ACC of GC is expected to be comparable to GDPS-SN.
Figure 13: Anomaly correlation coefficient against the ECMWF-HRES analyses with GDPS-CTL (blue) and GDPS-SN (red) for (a,d) northern hemisphere 500-hPa geopotential height, (b,e) tropical 850-hPa zonal wind, and (c,f) southern hemisphere 500-hPa geopotential height. Results are presented for 60 cases of boreal winter 2022 (left column: a–c) and 62 cases of boreal summer 2022 (right column: d–f). Red and blue markers denote statistically significant differences in favor of GDPS-SN and GDPS-CTL, respectively. No marker at a pressure level implies that the null hypothesis, stating that the statistics of the two samples are the same, cannot be rejected based on the 95th percentile. Grey arrows with printed numbers depict gains in predictability in terms of forecast hours.
@@ 6.4 Verification against surface observations
For verification at the surface level, combined SYNOP, METAR and Surface Weather and marine OBservations (SWOB; available only over Canada and distributed by ECCC) data were used for surface pressure, screen-level temperature, dewpoint depression and anemometer-level wind speed. For precipitation, ground observations of 24-hr accumulation, subject to collection and quality control by the CAnadian Precipitation Analysis system (CaPA; Lespinas et al. 2015), are used. Any data pertaining to stations with an altitude difference larger than 100 m with respect to GDPS orography, along with any observed wind speed below 1.5 $\text{m s}^{-1}$, are excluded.
@@
Figure 15: Q-Q plots of day-5 GC (black), GDPS-CTL (blue) and GDPS-SN (red) forecasts against combined SYNOP, SWOB and METAR for: (a, b) anemometer-level wind speeds and (c, d) screen-level temperatures over North America for (a, c) 30 cases of boreal winter valid at (a) 00 and (c) 12 UTC, and (b, d) 31 cases of boreal summer valid 00 UTC. Local standard time over North America is mostly between UTC-5 and UTC-8 h. Diamond symbols denote selected percentiles with their values printed. Light shadings indicate the 5%–95% confidence interval for the inverse cumulative distribution functions, based on the Kolmogorov-Smirnov statistic.
@@
The ability of the GPDS-SN to represent high-impact weather, such as high wind speeds and extreme temperatures, is not significantly different from the GDPS-CTL but, at the same time, it is substantially better than GC, which provides a strong example of the value added by the hybrid system over GC (Fig. 15). For example, the distribution of late-afternoon wind speeds over North America provided by GC shows a considerable shift toward lower values, which increases at high percentiles of the distribution, resulting in a -3 $\text{m s}^{-1}$ bias around ${}^{\text{th}}$ percentile (Figs. 15 a, b). GDPS-CTL also has a negative bias but affecting only high percentiles and, more importantly, GDPS-SN closely follows GDPS-CTL, with little or no sign of deceleration at extreme wind speeds for both seasons. GraphCast also suffers a 2-4 K cold bias at mid-to-high temperature percentiles over North America in summer, which has no impact on the GDPS-SN; its distribution aligns very well with the observations. GraphCast’s poor performance with respect to the extremes can be partly explained by its lack of well-resolved fine scales that are generally critical for representing the extremes. Furthermore, inconsistencies in the surface forcing between ERA5 and GDPS analyses may also be a contributing factor for the observed underestimation of temperature extremes by GC. As the spectral nudging configuration presented in this study only targets synoptic scales and the boundary layer is allowed to evolve freely, GDPS-SN is not adversely affected by GC’s limitations regarding the extremes.
@@
Spectral nudging also significantly improves boreal winter precipitation forecast guidance over North America for a large range of thresholds from 1 to 50 mm day-1, mostly by reducing the false alarm ratio, with no reduction of probability of detection, which results in a reduced moist frequency bias and an improved equitable threat score (Fig. 16). These improvements seem to be mostly generated via an improved prediction of synoptic-scale flow. There is, however, not much of an impact of spectral nudging on summer precipitation over North America (not shown). During summer, when weak synoptic forcings are prevalent, nudging only large scales is not expected to considerably improve precipitation forecasts, although extending nudging to the boundary layer may lead to some improvements by mitigating model biases.
@@
Figure 16: Quantitative precipitation forecast verification for day 5 assessed by 24-hr precipitation threshold (accumulations between 108 and 132 hr of integration) with (a) equitable threat score, (b) frequency bias, (c) probability of detection, and (d) false alarm ratio. Results from the 60 GDPS-CTL (blue) and GDPS-SN (red) cases of boreal winter are compared to ground observations over North America used by CaPA. Red and blue line markers denote statistically significant differences in favor of GDPS-SN and GDPS-CTL, respectively. Significance is computed by bootstrapping 3-day data blocks consisting of forecast-observation pairs from all stations of a given region. No marker at a threshold level implies that the null hypothesis, stating that the statistics of the two samples are the same, cannot be rejected based on the 90th percentile.
@@ 6.5 Tropical cyclone evaluation
Spectral nudging leads to an overall reduction in position error in the predicted TC trajectories (Fig. 17a). Notably, GDPS-CTL’s tendency of predicting storms that move too slowly is improved with GDPS-SN, especially at lead times beyond day 5 (Fig. 17b). There is also an indication of improvement with respect to GDPS-CTL’s tendency of predicting TCs that veer too much to the right from their observed tracks (Fig. 17c). Conversely, there is little or no significant impact of spectral nudging on the TC intensity, as measured by the minimum central pressure (not shown) and maximum sustained wind speed (Figs. 17d). Both GDPS-CTL and GDPS-SN have a similar weak-intensity bias, which is a well-known weakness of the GEM model (see McTaggart-Cowan et al. (2024), for more details). Severe TC weak-intensity biases are typical of state-of-the-art deterministic AI-based forecast models (e.g., Bouallègue et al. (2024)) due to the smoothing effect caused by learning to minimize the MSE. However, AI-based models, including GC, are generally capable of predicting the TC trajectories with enhanced accuracy (Lam et al., 2022). Therefore, the improvements in the predicted trajectories by GDPS-SN show that the proposed hybrid system is capable of leveraging the well-documented strength of the AI-based models, without sacrificing TC intensity. Furthermore, these results suggest that the weak-intensity bias for TCs could be reduced in future by focusing on improving the physics-based NWP component of the hybrid system.
Figure 17: Tropical cyclone tracking performance comparison of the GDPS-CTL (blue) and GDPS-SN (red) against IBTrACS data in three northern hemisphere basins (West Atlantic, East and West Pacific) for 62 cases of boreal summer 2022: (a) mean position error, (b) mean along-track position error, (c) mean cross-track position error, and (d) mean maximum wind speed error. Negative along-track errors imply cyclones move too slowly and positive cross-track errors imply cyclones veer to much to the right with respect to the observed trajectory. Red and blue markers denote statistically significant differences in favor of GDPS-SN and GDPS-CTL, respectively. No marker at a pressure level implies that the null hypothesis, stating that the statistics of the two samples are the same, cannot be rejected based on the 90th percentile. The number of cases for different forecast lead times are indicated in 17c.